If m and n are odd integers, which of the following must also be an odd integer?
a)(m+n) (m-n)
b)m(n+1)
c)mn+1
d)2m+n

Answer by Alan3354(69443)

If m and n are odd integers, which of the following must also be an odd integer?
a)(m+n) (m-n)
b)m(n+1)
c)mn+1
d)2m+n
-----------
Use m = 5 and n = 7:
a) 12*(-2) = -24
b) 5*6 = 30
c) 5*7+1 = 36
d) 2*5 + 7 = 17
------------
a, b & c are eliminated
Might be d
2*m is an even number.
Adding an odd number, n, --> an odd number.
It's d.